### v1

Abstract
Anatomical changes in head-and-neck cancer (HNC) patients during intensity-modulated proton therapy (IMPT) can shift the Bragg Peak of proton beams, risking tumor underdosing and organ-at-risk (OAR) overdosing. As a result, treatment replanning is often required to maintain clinically acceptable treatment quality. However, current manual replanning processes are often resource-intensive and time-consuming. In this work, we propose a patient-specific deep reinforcement learning (DRL) framework for automated IMPT replanning, with a reward-shaping mechanism based on a -point plan quality score designed to handle the competing clinical objectives in radiotherapy planning. We formulate the planning process as an RL problem where agents learn high-dimensional control policies to adjust plan optimization priorities to maximize plan quality. Unlike population-based approaches, our framework trains personalized agents for each patient using their planning CT (Computed Tomography) and augmented anatomies simulating anatomical changes (tumor progression and regression). This patient-specific approach leverages anatomical similarities along the treatment course, enabling effective plan adaptation. We implemented and compared two DRL algorithms, Deep Q-Network (DQN) and Proximal Policy Optimization (PPO), using dose-volume histograms (DVHs) as state representations and a -dimensional action space of priority adjustments. Evaluation on five HNC patients using actual replanning CT data showed that both DRL agents improved initial plan scores from to (DQN) and (PPO), surpassing the replans manually generated by a human planner (). Clinical validation confirms these improvements translate to better tumor coverage and OAR sparing across diverse anatomical changes. This work highlights the potential of DRL in addressing the geometric and dosimetric complexities of adaptive proton therapy, offering a promising solution for efficient offline adaptation and paving the way for online adaptive proton therapy.
1 Introduction
Intensity-modulated proton therapy (IMPT) provides highly conformal tumor coverage while sparing surrounding organs at risk (OARs), leveraging the unique dose deposition characteristics of the proton beam’s Bragg peak (Holliday et al., 2015; McKeever et al., 2016; Moreno et al., 2019). However, the Bragg peak makes IMPT highly sensitive to interfractional anatomical changes, such as tumor progression or regression, weight loss, and edema, that alter tissue density along the beam path (Huiskes et al., 2023; Sonke et al., 2019). These anatomical changes can shift the Bragg peak, degrading tumor coverage and/or OAR sparing, and often necessitate one or multiple manual replanning sessions during treatment to maintain clinical quality. Addressing these anatomical changes currently requires manual intervention: a multi-step process involving patient reimaging, anatomical recontouring, and iterative, trial-and-error plan optimization. This manual adaptation workflow is notoriously labor-intensive and typically takes several days, delaying optimal care and straining limited clinical resources (Kim et al., 2018; Leeman et al., 2017; Li et al., 2020).
From a machine learning (ML) perspective, automating treatment planning presents several challenges. First, the planning process involves navigating a high-dimensional optimization space with complex, non-linear relationships between planning parameters and resulting dose distributions (Burlacu et al., 2025; Wildman et al., 2024). Second, the multi-objective nature of treatment planning demands careful balancing of target coverage with the sparing of multiple OARs, each governed by its clinical constraints, thereby desiring methods that can effectively navigate trade-offs along a complex Pareto frontier (Barker Jr et al., 2004; Bobić et al., 2023; Sonke et al., 2019). Finally, the limited availability of patient data and substantial inter-patient variability in anatomy and tumor characteristics make it difficult to develop models that generalize well across large patient populations (Nikou et al., 2024; Visak et al., 2024; Volpe et al., 2021). These challenges are further amplified in head and neck cancer (HNC) treatment planning, which involves multiple treatment targets with varying prescription dose levels, and tumors often invade or abut several critical organs.
Reinforcement learning (RL) offers a compelling framework for tackling these sequential, multi-objective optimization challenges in HNC treatment planning. RL agents learn optimal strategies through interaction and reward feedback, making them particularly well-suited for navigating the complex, iterative process of balancing competing clinical objectives in treatment planning (Sutton et al., 1998). Specifically, RL holds the potential to automate the iterative decision-making process of priority adjustments in treatment planning (Eckardt et al., 2021; Ebrahimi and Lim, 2021; Moreau et al., 2021; Yang et al., 2024). Furthermore, RL’s inherent adaptability aligns with the vision of personalized online adaptive radiotherapy, where treatment plans can be dynamically adapted based on a patient’s tumor response and anatomical changes during treatment.
To realize this potential for automated IMPT replanning, we develop a patient-specific deep reinforcement learning (DRL) framework for the dosimetrically challenging HNC. Our approach formulates the priority tuning process during plan optimization as an RL problem, where the state is defined by dose–volume histograms (DVHs) for clinical target volumes (CTVs) and OARs, and the action space comprises predefined, clinically informed priority adjustments, enabling the agent to navigate the trade-offs inherent in this multi-objective optimization problem. Plan quality is quantified via a comprehensive -point scoring system that combines ProKnow standardized scoring criteria (Nelms et al., 2012) with institutional planning guidelines, and the reward is defined as the change in the plan quality score. Our framework trains personalized DRL agents using each patient’s anatomy captured by the initial planning CT (Computed Tomography) images and augmented anatomies simulating anatomical variations (tumor regression/progression). This patient-specific approach leverages the patient’s inherent anatomical consistency along the treatment course to optimize performance specifically for that individual rather than attempting to develop a one-size-fits-all solution.
Generalizable Insights about Machine Learning in the Context of Healthcare
Our work offers broader insights for applying machine learning in healthcare settings characterized by high variability and limited data:
Patient-specific architectures merit exploration in high-variability clinical applications. In medical settings characterized by limited data availability and substantial anatomical variability, developing ML models that generalize across large patient cohorts remains a fundamental challenge. However, despite undergoing anatomical changes, each patient typically maintains high anatomical consistency throughout the treatment course, which is a feature that has been entirely underutilized in traditional population-based ML models. By simulating plausible anatomical variations from a patient’s baseline anatomy to augment the limited patient-specific training datasets, our proposed patient-specific strategy offers a viable alternative to traditional population-based ML models and aligns closely with the goals of personalized precision medicine.
Clinically-informed reward design ensures alignment. Directly incorporating established clinical planning guidelines and quantitative plan scoring metrics into the RL reward function is crucial. This ensures that learned policies optimize for objectives recognized as clinically valid and relevant to real-world clinical decision-making priorities, fostering trust and clinical utility.
Structure | Planning Objective |
CTV1 (primary CTV) | |
CTV2 (secondary CTV) | |
CTV3 (tertiary CTV) | |
Brainstem (BRS) | |
Spinal Cord (SC) | |
Mandible (MAN) | |
Larynx (LAR) | |
Pharynx (PHY) | |
Left Parotid (PARL) | |
Right Parotid (PARR) | |
Left Cochlea (COCHL) | |
Right Cochlea (COCHR) | |
Left Submandibular Gland (SMGL) | |
Right Submandibular Gland (SMGR) | |
Esophagus (ESO) |
2 Related Work
RL has increasingly shown promise for automating radiotherapy treatment planning. Early demonstrations in cervical cancer high-dose-rate brachytherapy introduced a DRL-based Virtual Treatment Planner (VTP) that emulated human planners by observing DVH inputs and adjusting parameter weights, yielding plans that outperformed both the initial and human-generated plans (Shen et al., 2019). Building on this success, the framework was extended to external beam radiotherapy for prostate cancer (Shen et al., 2020), where separate DQN subnetworks adapted parameters in a relatively simple scenario with a single target and two OARs. Subsequent efforts by the same research group focused on improving training efficiency through rule-based adjustments informed by human-planner experience (Shen et al., 2021a) and scalability through a hierarchical VTP network that decomposed the planning process into structure selection, parameter selection, and action adjustment (Shen et al., 2021b), although still limited to single-target cases with few OARs. Diverging from neural network-based approaches, Zhang et al. (2020) developed an interpretable planning bot for pancreatic stereotactic body radiation therapy using linear function approximation for Q-learning. However, these approaches were primarily designed for anatomically simpler treatment sites and might struggle with sites like HNC, characterized by complex inter-structure trade-off relationships.
HNC represents a substantially more challenging planning context due to multiple target volumes with varying prescription doses and numerous critical structures in close proximity, demanding tight dosimetric trade-offs even for experienced planners. Recent RL adaptations to HNC planning have attempted to navigate this high-dimensional optimization space with varied strategies. Gao et al. (2024) adapted the hierarchical VTP approach for HNC photon therapy (whereas this work focuses on proton therapy), employing two DQN subnetworks for parameter selection and adjustment direction determination while representing states as plan quality scores rather than DVHs. To manage the high dimensionality arising in HNC planning, they manually fixed most of the potential planning parameters in their case based on dosimetrist expertise, allowing their VTP to learn adjustments for only unique priority parameters. Similarly, grappling with the high-dimensional action space, Wang and Chang (2024) explored policy-gradient methods with a transformer-based PPO agent for continuous parameter adjustments. Their approach proposed dynamically reducing complexity by adjusting parameters only for the subset of structures currently exhibiting the lowest quality scores. However, this strategy may not accommodate cases where certain OARs must be strategically compromised to maintain target coverage—a common requirement in complex HNC planning scenarios.
Notably, all the aforementioned approaches rely on population-based models trained across diverse patient cohorts. While effective at capturing common anatomical patterns, such models face challenges adapting to patient-specific anatomical changes that necessitate treatment replanning during HNC’s fractionated radiotherapy (Caudell et al., 2017; Ma et al., 2022; Maniscalco et al., 2023; Visak et al., 2024). This limitation, stemming from the under-utilization of intra-patient anatomical consistency, restricts their ability to generate optimally adapted plans tailored to each patient’s unique progression. Motivated by these challenges, we introduce a novel patient-specific RL framework for IMPT replanning in HNC. By developing and optimizing the planning policy to each individual’s unique anatomy and dosimetric constraints from the outset, our approach inherently leverages intra-patient consistency patterns. This personalization is crucial for effectively managing the high-dimensional optimization space and is potentially better suited to addressing the adaptive replanning demands inherent in HNC treatment, overcoming key limitations of existing population-based strategies.
3 Methods
3.1 Treatment Plan Optimization
In treatment planning systems (TPS), the patient anatomy (CTVs and OARs) is discretized into voxels. The dose deposited in each voxel is linearly dependent on adjustable beamlet intensities, represented by the fluence vector . This relationship is pre-calculated and stored as a dose influence matrix , where each element quantifies the dose contribution from unit intensity of beamlet to voxel (Gorissen, 2022):
The goal of treatment plan optimization is to determine the optimal beamlet intensity vector that minimizes dose deviations in CTV voxels from their prescribed dose while reducing the dose to OARs, which can be formulated as:
Here, is the multi-objective treatment planning loss function, where is the vector of beamlet intensities to be optimized. Each component represents the intensity or weight of the -th beamlet. The constraint explicitly ensures that all beamlet intensities are non-negative. The formulation incorporates CTVs, where indexes each CTV. Similarly, OARs are considered, where indexes each OAR. For the patient cohort in this study, and . The weighting parameters and serve as treatment planning priorities (TPPs) that balance competing clinical objectives: achieving prescribed dose coverage in the -th CTV while minimizing dose exposure to the -th OAR. These weights are the primary optimization variables manipulated by our DRL agents. and denote the dose influence matrices for the -th CTV and -th OAR. is the prescribed dose vector for the -th CTV.
In practice, achieving a perfect homogeneous dose distribution within CTVs is not feasible; hot and cold spots are often unavoidable. Similarly, since radiation beams must often pass through some OARs to reach the CTVs, complete OAR sparing from radiation dose is unrealistic. Importantly, different OARs have varying dose tolerances based on their biological characteristics; as long as the dose to each OAR remains within clinically acceptable thresholds, the risk of severe toxicity is minimized. Hence, to reflect these clinical realities and achieve realistic plan optimization, in this study we implemented a DVH-based cost function incorporating multiple planning objectives. The planning objectives involved in our IMPT plan optimization for HNC are listed in Table 1, where represents the volume of a structure receiving a dose equal to or greater than , while denotes the minimum dose received by the hottest volume of a structure. The trade-offs between these competing objectives are controlled by objective-specific priorities (), which are traditionally set manually through iterative tuning. The DVH-based cost function (Eqn. 2) was optimized using an in-house dose calculation engine, which serves as the environment for our RL-based planning optimization.
3.2 Reinforcement Learning for Automatic Replanning
Converting the aforementioned priority tuning during inverse plan optimization into an RL problem offers a solution to the time-consuming, labor-intensive manual adjustment process. We model the priority tuning as a finite Markov Decision Process defined by the tuple (), where:
State Space (): Each state represents the DVHs of the CTVs and typical OARs (Table 1), represented as a normalized matrix where corresponds to anatomical structures and to discretized dose-volume bins.
Action Space (): A discrete set of clinically constrained actions representing priority adjustments for different planning objectives, designed based on clinical experience.
Transition Dynamics (): The transition model is governed by a treatment optimization engine, with capturing how priority adjustment actions affect the resulting dose distribution.
Reward Function (): The reward function defines the immediate reward obtained after taking action in state . It is defined as , where is the plan quality score of the intermediate plan dose, generated by taking action . We designed a -point scoring system based on the standardized ProKnow scoring criteria (Nelms et al., 2012) and institution-specific planning guidelines. It includes a set of dosimetric metrics for the structures of interest. The final score was calculated as the sum of the scores across all these metrics, with higher scores indicating better plan quality.
Discount Factor (): balances immediate dosimetric improvements with long-term overall plan quality.
The goal is to learn an optimal policy that dictates the sequence of priority adjustments leading to the highest possible cumulative reward, thereby yielding high-quality treatment plans:
where represents a policy that maps DVH states to priority adjustment actions, and is the planning horizon. The expectation is taken over the stochasticity in state transitions under policy starting from initial state .
The optimal policy formulation provides a theoretical objective, but solving for directly is challenging in high-dimensional state spaces like those encountered in IMPT replanning. To address this challenge, we implemented two state-of-the-art DRL algorithms that have demonstrated success in complex sequential decision-making tasks.
Deep Q-Network (DQN)
In Q-learning, the objective is to learn the optimal action-value function , which represents the maximum expected return achievable by taking action in state and following the optimal policy thereafter (Sutton et al., 1998; Watkins and Dayan, 1992). For proton therapy planning, this corresponds to predicting which priority adjustment () will yield the greatest long-term improvement in plan quality. The optimal Q-function satisfies the Bellman equation:
DQN approximates this optimal Q-function using a deep neural network with weights . The network is trained by minimizing a loss function that measures the temporal difference (TD) error between the predicted Q-value and the target Q-value (Mnih et al., 2015). The loss function at iteration is given by:
where is a replay buffer storing past experiences (often with prioritized sampling based on TD error), and the target is calculated using a separate target network with delayed weights , typically updated periodically:
The DQN agent learns a policy by selecting actions with the highest Q-value for a given state (e.g., using an -greedy strategy to balance exploration and exploitation).
Proximal Policy Optimization (PPO)
While DQN learns Q-values of state-action pairs and derives policies through action selection mechanisms, PPO takes a more direct approach. PPO is an actor-critic policy gradient algorithm that explicitly learns a policy parameterized by weights and a value function parameterized by weights . The policy (actor) determines actions based on the current state, while the value function (critic) approximates the value of the current state, providing a baseline for evaluating the actor’s performance. Policy gradient methods update the actor’s parameters by following the gradient of an objective function that aims to maximize the expected return (Schulman et al., 2017). To ensure stable training, PPO employs a clipped surrogate objective function that prevents excessively large policy updates by maintaining the new policy within a trust region of the old policy. The objective function for the actor is defined as:
where is the policy from the previous iteration, is the current policy, is a hyperparameter controlling the clipping range, and is the generalized advantage estimate (GAE) (Schulman et al., 2016):
where is the discount factor, controls the bias-variance tradeoff, and represents the value function approximation. GAE estimates how much better an action is compared to the average action in a given state.
Simultaneously, the critic is trained to accurately estimate the state value by minimizing the mean squared error between its predictions and the actual discounted return:
where is the discounted return and is the experience replay of the current policy.
In essence, PPO aims to find optimal policy and value function parameters by iteratively minimizing a combined objective function that includes both the actor loss and the critic loss, thereby approximately solving the minimization problem given by:
where and are coefficients controlling the relative importance of the value function loss and the entropy bonus , respectively. The entropy bonus encourages exploration.
4 Patient Cohort
4.1 Cohort Selection
We validated our approach through a retrospective study involving five HNC patients treated with IMPT at Hospital A. All patients underwent replanning during their treatment course due to substantial anatomical changes. A CT scan for treatment planning (referred to as pCT) was acquired weeks before the treatment course, and a second CT scan was obtained for replanning (rpCT) during the course when substantial anatomical changes were observed and warranted replanning. These five cases were deliberately selected for their varied tumor characteristics and spatial relationships to surrounding OARs, and all cases involved three CTVs, representing challenging scenarios for treatment planning. The primary CTV (CTV1) for all cases was prescribed to receive over fractions, while the prescription doses for the secondary (CTV2) and tertiary (CTV3) targets varied among cases, as detailed in Table 2. This variability in prescription doses poses a significant challenge for automated treatment planning and replanning in HNC, as population-based models often struggle to generalize across diverse clinical objectives, let alone the substantial variation in patient anatomy and tumor characteristics.
Case ID | Volume on pCT () | Volume on rpCT () | No. of Fractions | Prescription (Gy RBE) | ||||||
CTV1 | CTV2 | CTV3 | CTV1 | CTV2 | CTV3 | CTV1 | CTV2 | CTV3 | ||
P1 | 35 | |||||||||
P2 | 35 | |||||||||
P3 | 35 | |||||||||
P4 | 35 | |||||||||
P5 | 35 |
4.2 Data Extraction
Relevant patient data, including the pCT and rpCT, and their corresponding CTV and OAR contours, were retrieved from an internal database in DICOM format. All the contours were manually delineated by attending physicians for treatment planning or replanning of the actual treatments. These DICOM files of CT images and contours were imported into the open-source treatment planning toolkit matRad (Wieser et al., 2017) to precalculate dose influence matrices () required for plan optimization, using the same beam arrangements (e.g., isocenter location, number of beams, and beam angles) as those used in the actual treatments. The dose matrix calculated from the pCT and its associated contours was used for training the DRL agents, while the matrix calculated from the rpCT and corresponding contours was used for testing.
To train a patient-specific agent capable of adjusting the priorities among planning objectives to adapt the treatment plan to anatomical changes, we further generated augmented anatomies by simulating anatomical changes. In this proof-of-concept study, we simulated only two variation scenarios, i.e., tumor progression and regression, by expanding the original CTV contours on the pCT by or shrinking them by , respectively. These variations not only changed the CTV volumes but also altered their spatial relationships with surrounding OARs, increasing the diversity of the patient-specific training dataset in terms of anatomy and planning complexity. The pCT images, along with each set of modified contours, were also imported into matRad to calculate the corresponding dose influence matrices for the simulated anatomy variations.
5 Experiments & Results
In this section, we present the results of our experiments and evaluate the effectiveness of RL-based planners for automated IMPT replanning in HNC treatment.
Case ID | Initial plan | Manual plan | DQN plan | PPO plan |
P1 | 122.56 | 130.85 | 131.05 | 137.88 |
P2 | 83.14 | 132.09 | 136.20 | 138.05 |
P3 | 132.35 | 140.44 | 147.25 | 147.52 |
P4 | 132.14 | 143.94 | 146.10 | 148.77 |
P5 | 132.96 | 138.67 | 138.32 | 141.50 |
mean std |
5.1 Experimental Setup
For each patient, both DQN and PPO agents were trained using the patient’s original anatomy from the pCT and two augmented anatomical variations. The state representation included the normalized DVH curves for all structures (dim=), while the discrete action space consisted of priority adjustments for CTVs and OARs. Each agent was trained for episodes, with each episode consisting of up to priority adjustment steps or terminating early if the maximum plan quality score of was achieved. Algorithm 1 provides a detailed overview of this DRL-based replanning process, illustrating the integration of priority tuning and plan optimization within our experimental framework.
The trained agents were then tested on the patient’s new anatomy captured in the rpCT. Specifically, starting from an initial default priority set , the agents performed priority tuning within tuning steps, where each step involved selecting a tuning action based on the plan quality of the intermediate plan and was followed by inverse plan optimization. For comparison, a new treatment plan was also manually generated for each case on the new anatomy by a human planner through manual priority tuning. The performance of the agents was assessed by comparing the plan quality scores and dosimetric metrics of the agent-generated plans with those of the manually created plans.
To facilitate the interaction between the RL agents and the treatment planning process, we developed a custom OpenAI Gymnasium-compatible environment that simulates the treatment planning workflow. At each timestep, the environment processes the current beamlet weights along with the dose influence matrices and for each CTV and OAR . It then solves the DVH-based optimization problem using the priority set , modified by the agent-selected action , updates , and generates a new dose distribution , which is subsequently used to compute the plan score and reward via Eqn. 1. Each action updates priority weights according to:
where ensures weights remain within valid bounds. All experiments were conducted on a workstation equipped with an NVIDIA RTX Ada Generation GPU with of memory.
5.2 Plan Quality Comparisons
The same -point scoring system used to calculate rewards during DRL training was employed to quantify the plan quality of treatment plans generated for each patient’s new anatomy captured in the rpCT. The resulting plan quality scores are presented in Table 3. The score for each dosimetric metric of the generated plans are summarized in Appendix A. As shown, both DQN and PPO agents effectively adjusted the priorities of planning objectives, significantly improving the plan quality compared to the initial plans generated using the default priority set. Notably, for every case, both agents achieved plans that were comparable to or better than those manually created by a human planner. PPO-generated plans consistently achieved the highest quality scores across all cases (mean: ), outperforming both manual plans (, via paired t-test) and DQN-generated plans (). Furthermore, the PPO-generated plans exhibited lower score variability (, vs. for manual plans and for DQN-generated plans), indicating enhanced consistency and reliability in adapting to anatomical changes.
5.3 Dosimetric Analysis
The detailed dosimetric metrics of each structure for the manually generated plans and those generated by the DQN and PPO agents are presented in Table 4. Clinically acceptable target dose coverage was achieved in all the three plans for every case, with the coverage of primary CTV () no less than . In terms of dose homogeneity inside the primary CTV (CTV1), both RL agents reduced hot spots in three patients while maintaining excellent coverage. Most notably, in patient P4, DQN and PPO reduced the maximum dose within CTV1 by and , respectively, compared to the manual plan. PPO demonstrated superior OAR sparing in most patients, reducing mean dose to the larynx (LAR) ( (PPO) vs. (Manual)) and to the bilateral parotids (PARL and PARR) ( (PPO) vs. (Manual)), compared to manual replanning. The effectiveness of PPO was particularly evident in case P5, where it slightly compromised the dose coverage of the tertiary CTV (CTV3, ) to substantially reduce the dose to the right submandibular gland (SMGR), bringing its mean dose closer to the dose tolerance ( ) and resulting in a significantly higher plan score. In contrast, the manual and DQN-generated plans for this case substantially exceeded the dose tolerance of the SMGR, with mean doses of and , respectively. Figure 1 compares the DVH of the plans generated manually by a human planner against the automated plans generated using RL agents (DQN and PPO) for patients P4 and P5. DVH curves and dosimetric results for all cases are provided in Appendix A.
Structure | Metric | P1 | P2 | P3 | P4 | P5 | ||||||||||
M | Q | P | M | Q | P | M | Q | P | M | Q | P | M | Q | P | ||
CTV1 | 97.99 | 98.01 | 97.99 | 97.99 | 98.01 | 97.99 | 97.99 | 98.00 | 97.99 | 98.01 | 97.97 | 98.01 | 97.99 | 97.99 | 98.00 | |
81.86 | 81.02 | 82.24 | 82.30 | 83.03 | 83.61 | 81.52 | 81.00 | 79.68 | 81.23 | 77.82 | 78.58 | 77.26 | 80.82 | 77.82 | ||
CTV2 | 97.97 | 97.99 | 98.56 | 97.94 | 98.27 | 98.04 | 98.39 | 98.39 | 98.29 | 98.55 | 98.14 | 98.19 | 99.51 | 98.50 | 99.19 | |
CTV3 | 97.84 | 98.03 | 99.00 | 97.76 | 98.27 | 98.18 | 98.18 | 98.33 | 98.02 | 98.02 | 97.94 | 98.03 | 99.12 | 97.98 | 97.53 | |
BRS | 23.29 | 24.01 | 24.17 | 19.59 | 19.58 | 20.04 | 10.69 | 10.55 | 10.77 | 2.87 | 2.82 | 2.85 | 13.69 | 12.48 | 14.07 | |
SC | 37.63 | 37.67 | 40.60 | 22.63 | 22.70 | 22.85 | 31.81 | 32.35 | 32.17 | 13.85 | 14.04 | 14.98 | 30.58 | 30.23 | 32.06 | |
MAN | 1.73 | 0.58 | 2.18 | 1.29 | 1.07 | 1.51 | 0.13 | 0.00 | 1.28 | 0.22 | 0.09 | 0.09 | 0.08 | 0.67 | 0.20 | |
LAR | 43.76 | 43.84 | 40.15 | 36.51 | 36.48 | 36.76 | 34.76 | 34.84 | 34.91 | 45.29 | 44.51 | 39.29 | 32.10 | 32.74 | 30.86 | |
PHY | 0.00 | 0.00 | 0.00 | 49.13 | 48.95 | 49.41 | 0.00 | 0.00 | 0.00 | 44.00 | 43.17 | 42.13 | 37.73 | 38.98 | 35.12 | |
PARL | 14.43 | 14.57 | 15.56 | 26.00 | 26.00 | 22.74 | 20.39 | 20.19 | 20.52 | 21.76 | 21.79 | 21.86 | 10.04 | 9.63 | 10.12 | |
PARR | 34.60 | 34.50 | 23.86 | 28.41 | 25.15 | 25.81 | 32.68 | 25.97 | 25.71 | 10.08 | 10.28 | 10.19 | 21.56 | 21.81 | 21.69 | |
COCHL | 4.03 | 4.15 | 3.96 | 39.34 | 39.82 | 33.55 | 0.02 | 0.02 | 0.02 | 6.92 | 6.93 | 6.97 | 0.00 | 0.00 | 0.00 | |
COCHR | 2.09 | 2.07 | 2.42 | 8.44 | 8.47 | 8.56 | 0.84 | 0.88 | 0.84 | 0.02 | 0.02 | 0.02 | 0.00 | 0.00 | 0.00 | |
SMGL | 59.73 | 60.59 | 60.07 | 72.57 | 72.11 | 72.68 | 0.00 | 0.00 | 0.00 | 0.00 | 0.00 | 0.00 | 68.14 | 71.57 | 68.15 | |
SMGR | 0.00 | 0.00 | 0.00 | 72.83 | 72.51 | 73.06 | 0.00 | 0.00 | 0.00 | 36.87 | 37.30 | 31.98 | 67.28 | 70.89 | 37.60 | |
ESO | 11.09 | 11.09 | 11.57 | 6.23 | 6.26 | 6.24 | 13.26 | 13.35 | 13.31 | 1.36 | 1.38 | 1.38 | 6.41 | 5.85 | 6.69 | |
Plan Score (max=150): | 130.85 | 131.05 | 137.88 | 132.09 | 136.20 | 138.05 | 140.44 | 147.25 | 147.52 | 143.94 | 146.10 | 148.77 | 138.67 | 138.32 | 141.50 |
Note: OAR abbreviations - BRS: Brainstem, SC: Spinal Cord, MAN: Mandible, LAR: Larynx, PHY: Pharynx, PARL/PARR: Left/Right Parotid, COCHL/COCHR: Left/Right Cochlea, SMLG/SMGR: Left/Right Submandibular Gland, ESO: Esophagus.
6 Discussion
Our study demonstrates that RL-based automated IMPT replanning, particularly using PPO, consistently generates treatment plans of higher quality for HNC patients experiencing anatomical changes during the treatment course, compared to those manually generated by a human planner. These improvements have several important implications for adaptive proton therapy workflows.
6.1 Clinical Significance
The marked improvement in plan quality scores across all patients indicates that RL-based automated replanning can standardize high-quality treatment planning. This may lead to reduced treatment-related toxicities while maintaining effective tumor control. For example, the reduction in parotid gland dose has been correlated with decreased risk of xerostomia, with studies suggesting that each reduction in mean dose translates to approximately reduction in xerostomia risk (Castelli et al., 2015; Chao et al., 2001). As shown in Table 4, patients P1 and P2 benefited from substantial parotid sparing (reductions of and , respectively), while patients P1 and P4 experienced notable laryngeal sparing (reductions of and ). Importantly, these dosimetric improvements were achieved without compromising target coverage, underscoring the ability of RL methods to address the geometric and dosimetric complexities in HNC treatment planning and balance complex trade-offs to achieve clinically acceptable treatment plans. The proposed RL-based automated replanning approach offers a promising solution for efficient offline adaptation and paves the way for online adaptive proton therapy.
6.2 RL-based Planning and Patient-specific Benefits
Both RL algorithms outperformed manual replanning, with PPO consistently delivering superior plans compared to DQN despite operating within the same discrete action space. This finding aligns with recent literature suggesting policy-based methods like PPO effectively navigate the complex reward landscapes common in radiotherapy planning (Wang and Chang, 2024). PPO’s strategy of performing small, incremental parameter updates helps prevent detrimental large policy shifts, a feature particularly well-suited to the radiotherapy optimization problem.
Analysis of patient-specific outcomes relative to anatomical features (Tables 2 and 3) underscores this advantage. PPO consistently delivered the highest plan scores, outperforming manual replans across the cohort irrespective of initial target volumes (rpCT CTV1 range: ) or the extent of their volumetric changes between pCT and rpCT (e.g., P2 CTV3: expansion). Notably, it achieved the highest absolute score () for the patient with the smallest CTV1/CTV2 (P4), while providing clinically meaningful gains over manual plans for large-volume cases (PPO plan score vs. manual for P3: ; P5: ). Overall, the results strongly suggest that DRL-based automation offers significant dosimetric advantages for IMPT replanning. By consistently optimizing plans towards high scores, the DRL framework holds considerable promise for improving plan quality for HNC patients undergoing significant anatomical changes.
6.3 Limitations and Future Directions
While our findings demonstrate the feasibility of deep reinforcement learning (DRL) for adaptive IMPT replanning, important limitations should be acknowledged alongside opportunities for future work. First, the retrospective evaluation was conducted on a limited cohort of five patients. However, this cohort is representative of challenging HNC cases with three CTVs of different prescription levels and common anatomical change patterns. Future work should apply our patient-specific approach to a larger, more diverse cohort of patients with varying CTV configurations, rigorously assessing the framework’s adaptability across different anatomical presentations. The demonstrated superiority of our patient-specific DRL agents over manual plans further motivates comprehensive evaluation in prospective studies.
Second, the current framework’s reliance on simulated anatomical changes, while methodologically necessary for controlled RL development, necessitates validation with real longitudinal anatomical changes observed during actual treatment courses. Simulated target volume expansions and contractions may not fully capture the complex, heterogeneous tissue deformations, weight loss patterns, tumor shrinkage dynamics, and normal tissue responses that occur during radiotherapy. Future work should incorporate actual or additional anatomical change scenarios for training to enhance the performance of the DRL agents further.
Finally, methodological advancements to the RL agent itself could yield further dosimetric improvements. While our current approach focuses on clinically interpretable discrete priority adjustments, exploring continuous action spaces could potentially allow for finer control over planning parameters and enable more nuanced treatment planning. Exploring more sophisticated state representations may further reduce computational demands and enhance plan quality. To address the computational overhead associated with patient-specific training, transfer learning techniques might be needed. Such techniques could facilitate the development of readily deployable models, requiring only minimal fine-tuning for new patients.
7 Conclusion
Reinforcement learning, particularly PPO-based approaches, offers a compelling approach to automated replanning in HNC IMPT. The patient-specific nature of our RL framework enables tailored optimization strategies that adapt to the unique anatomical and dosimetric challenges of each patient. Our findings demonstrate a consistent generation of superior treatment plans compared to manual planning, potentially reducing planning time and improving plan quality. These results suggest that RL-based solutions can significantly enhance IMPT workflows, ultimately benefiting cancer patients through reduced toxicities and effective tumor control.
This project is supported by the National Institute of Health under Award Numbers
R01DE033512 and R37CA272755.
Appendix A Complete Dosimetric Results
Delineated | Dosimetric Endpoint | Manual | DQN | PPO | ||||
Structure | Metric | Score | Value | Score | Value | Score | Value | Score |
CTV1 | 15 | 97.99 | 14.99 | 98.01 | 15.00 | 97.99 | 14.99 | |
5 | 81.86 | 3.31 | 81.02 | 3.43 | 82.24 | 3.25 | ||
CTV2 | 15 | 97.97 | 14.98 | 97.99 | 14.99 | 98.56 | 15.00 | |
CTV3 | 15 | 97.84 | 14.89 | 98.03 | 15.00 | 99.00 | 15.00 | |
BRS | 21 | 23.29 | 21.00 | 24.01 | 21.00 | 24.17 | 21.00 | |
SC | 21 | 37.63 | 17.18 | 37.67 | 17.17 | 40.60 | 15.70 | |
MAN | 5 | 1.73 | 5.00 | 0.58 | 5.00 | 2.18 | 5.00 | |
LAR | 7 | 43.76 | 5.50 | 43.84 | 5.46 | 40.15 | 6.94 | |
PHY | 7 | 0.00 | 7.00 | 0.00 | 7.00 | 0.00 | 7.00 | |
PARL | 7 | 14.43 | 7.00 | 14.57 | 7.00 | 15.56 | 7.00 | |
PARR | 7 | 34.60 | 0.00 | 34.50 | 0.00 | 23.86 | 7.00 | |
COCHL | 5 | 4.03 | 5.00 | 4.15 | 5.00 | 3.96 | 5.00 | |
COCHR | 5 | 2.09 | 5.00 | 2.07 | 5.00 | 2.42 | 5.00 | |
SMGL | 5 | 59.73 | 0 | 60.59 | 0 | 60.07 | 0 | |
SMGR | 5 | 0.00 | 5.00 | 0.00 | 5.00 | 0.00 | 5.00 | |
ESO | 5 | 11.09 | 5.00 | 11.09 | 5.00 | 11.57 | 5.00 | |
Cumulative Total | 150 | – | 130.85 | – | 131.05 | – | 137.88 |
Delineated | Dosimetric Endpoint | Manual | DQN | PPO | ||||
Structure | Metric | Score | Value | Score | Value | Score | Value | Score |
CTV1 | 15 | 97.99 | 14.99 | 98.01 | 15.00 | 97.99 | 14.99 | |
5 | 82.30 | 3.24 | 83.03 | 3.14 | 83.61 | 3.06 | ||
CTV2 | 15 | 97.94 | 14.96 | 98.27 | 15.00 | 98.04 | 15.00 | |
CTV3 | 15 | 97.76 | 14.84 | 98.27 | 15.00 | 98.18 | 15.00 | |
BRS | 21 | 19.59 | 21.00 | 19.58 | 21.00 | 20.04 | 21.00 | |
SC | 21 | 22.63 | 21.00 | 22.70 | 21.00 | 22.85 | 21.00 | |
MAN | 5 | 1.29 | 5.00 | 1.07 | 5.00 | 1.51 | 5.00 | |
LAR | 7 | 36.51 | 7.00 | 36.48 | 7.00 | 36.76 | 7.00 | |
PHY | 7 | 49.13 | 7.00 | 48.95 | 7.00 | 49.41 | 7.00 | |
PARL | 7 | 26.00 | 7.00 | 26.00 | 7.00 | 22.74 | 7.00 | |
PARR | 7 | 28.41 | 2.78 | 25.15 | 7.00 | 25.81 | 7.00 | |
COCHL | 5 | 39.34 | 3.27 | 39.82 | 3.07 | 33.55 | 5.00 | |
COCHR | 5 | 8.44 | 5.00 | 8.47 | 5.00 | 8.56 | 5.00 | |
SMGL | 5 | 72.57 | 0 | 72.11 | 0 | 72.68 | 0 | |
SMGR | 5 | 72.83 | 0 | 72.51 | 0 | 73.06 | 0 | |
ESO | 5 | 6.23 | 5.00 | 6.26 | 5.00 | 6.24 | 5.00 | |
Cumulative Total | 150 | – | 132.09 | – | 136.20 | – | 138.05 |
Delineated | Dosimetric Endpoint | Manual | DQN | PPO | ||||
Structure | Metric | Score | Value | Score | Value | Score | Value | Score |
CTV1 | 15 | 97.99 | 14.99 | 98.00 | 15.00 | 97.99 | 14.99 | |
5 | 81.52 | 3.35 | 81.00 | 3.43 | 79.68 | 3.62 | ||
CTV2 | 15 | 98.39 | 15.00 | 98.39 | 15.00 | 98.29 | 15.00 | |
CTV3 | 15 | 98.18 | 15.00 | 98.33 | 15.00 | 98.02 | 15.00 | |
BRS | 21 | 10.69 | 21.00 | 10.55 | 21.00 | 10.77 | 21.00 | |
SC | 21 | 31.81 | 20.09 | 32.35 | 19.82 | 32.17 | 19.91 | |
MAN | 5 | 0.13 | 5.00 | 0.00 | 5.00 | 1.28 | 5.00 | |
LAR | 7 | 34.76 | 7.00 | 34.84 | 7.00 | 34.91 | 7.00 | |
PHY | 7 | 0.00 | 7.00 | 0.00 | 7.00 | 0.00 | 7.00 | |
PARL | 7 | 20.39 | 7.00 | 20.19 | 7.00 | 20.52 | 7.00 | |
PARR | 7 | 32.68 | 0.00 | 25.97 | 7.00 | 25.71 | 7.00 | |
COCHL | 5 | 0.02 | 5.00 | 0.02 | 5.00 | 0.02 | 5.00 | |
COCHR | 5 | 0.84 | 5.00 | 0.88 | 5.00 | 0.84 | 5.00 | |
SMGL | 5 | 0.00 | 5.00 | 0.00 | 5.00 | 0.00 | 5.00 | |
SMGR | 5 | 0.00 | 5.00 | 0.00 | 5.00 | 0.00 | 5.00 | |
ESO | 5 | 13.26 | 5.00 | 13.35 | 5.00 | 13.31 | 5.00 | |
Cumulative Total | 150 | – | 140.44 | – | 147.25 | – | 147.52 |
Delineated | Dosimetric Endpoint | Manual | DQN | PPO | ||||
Structure | Metric | Score | Value | Score | Value | Score | Value | Score |
CTV1 | 15 | 98.01 | 15.00 | 97.97 | 14.98 | 98.01 | 15.00 | |
5 | 81.23 | 3.40 | 77.82 | 3.88 | 78.58 | 3.77 | ||
CTV2 | 15 | 98.55 | 15.00 | 98.14 | 15.00 | 98.19 | 15.00 | |
CTV3 | 15 | 98.02 | 15.00 | 97.94 | 14.96 | 98.03 | 15.00 | |
BRS | 21 | 2.87 | 21.00 | 2.82 | 21.00 | 20.04 | 21.00 | |
SC | 21 | 13.85 | 21.00 | 14.04 | 21.00 | 22.85 | 21.00 | |
MAN | 5 | 0.22 | 5.00 | 0.09 | 5.00 | 0.09 | 5.00 | |
LAR | 7 | 45.29 | 3.30 | 44.51 | 5.20 | 39.29 | 7.00 | |
PHY | 7 | 44.00 | 7.00 | 43.17 | 7.00 | 42.13 | 7.00 | |
PARL | 7 | 21.76 | 7.00 | 21.79 | 7.00 | 21.86 | 7.00 | |
PARR | 7 | 10.08 | 7.00 | 10.28 | 7.00 | 10.19 | 7.00 | |
COCHL | 5 | 6.92 | 5.00 | 6.93 | 5.00 | 6.97 | 5.00 | |
COCHR | 5 | 0.02 | 5.00 | 0.02 | 5.00 | 0.02 | 5.00 | |
SMGL | 5 | 0.00 | 5.00 | 0.00 | 5.00 | 0.00 | 5.00 | |
SMGR | 5 | 36.87 | 4.25 | 37.30 | 4.08 | 37.98 | 5.00 | |
ESO | 5 | 1.36 | 5.00 | 1.38 | 5.00 | 1.38 | 5.00 | |
Cumulative Total | 150 | – | 143.94 | – | 146.10 | – | 148.77 |
Delineated | Dosimetric Endpoint | Manual | DQN | PPO | ||||
Structure | Metric | Score | Value | Score | Value | Score | Value | Score |
CTV1 | 15 | 97.99 | 14.99 | 97.99 | 14.99 | 98.00 | 15.00 | |
5 | 77.26 | 3.96 | 80.82 | 3.45 | 77.82 | 3.88 | ||
CTV2 | 15 | 99.51 | 15.00 | 98.50 | 15.00 | 99.19 | 15.00 | |
CTV3 | 15 | 99.12 | 15.00 | 97.98 | 14.99 | 97.53 | 14.69 | |
BRS | 21 | 13.69 | 21.00 | 12.48 | 21.00 | 14.07 | 21.00 | |
SC | 21 | 30.58 | 20.71 | 30.23 | 20.88 | 32.06 | 19.97 | |
MAN | 5 | 0.08 | 5.00 | 0.67 | 5.00 | 0.20 | 5.00 | |
LAR | 7 | 32.10 | 7.00 | 32.74 | 7.00 | 30.86 | 7.00 | |
PHY | 7 | 37.73 | 7.00 | 38.98 | 7.00 | 35.12 | 7.00 | |
PARL | 7 | 10.04 | 7.00 | 9.63 | 7.00 | 10.12 | 7.00 | |
PARR | 7 | 21.56 | 7.00 | 21.81 | 7.00 | 21.69 | 7.00 | |
COCHL | 5 | 0.00 | 5.00 | 0.00 | 5.00 | 0.00 | 5.00 | |
COCHR | 5 | 0.00 | 5.00 | 0.00 | 5.00 | 0.00 | 5.00 | |
SMGL | 5 | 68.14 | 0 | 71.57 | 0 | 68.15 | 0 | |
SMGR | 5 | 67.28 | 0 | 70.89 | 0 | 37.60 | 0 | |
ESO | 5 | 6.41 | 5.00 | 5.85 | 5.00 | 6.69 | 5.00 | |
Cumulative Total | 150 | – | 138.67 | – | 138.32 | – | 141.50 |

### v2

@@
 As a result, treatment replanning is often required to maintain clinically acceptable treatment quality.
-However, current manual replanning processes are often resource-intensive and time-consuming.
-In this work, we propose a patient-specific deep reinforcement learning (DRL) framework for automated IMPT replanning, with a reward-shaping mechanism based on a -point plan quality score designed to handle the competing clinical objectives in radiotherapy planning.
-We formulate the planning process as an RL problem where agents learn high-dimensional control policies to adjust plan optimization priorities to maximize plan quality.
-Unlike population-based approaches, our framework trains personalized agents for each patient using their planning CT (Computed Tomography) and augmented anatomies simulating anatomical changes (tumor progression and regression).
+However, current manual replanning processes are often resource intensive and time consuming.
+In this work, we propose a patient-specific deep reinforcement learning (DRL) framework for automated IMPT replanning, with a reward-shaping mechanism based on a -point plan quality score designed to handle competing clinical objectives in radiotherapy planning.
+We formulate the planning process as a reinforcement learning (RL) problem where agents learn high-dimensional control policies to adjust plan optimization priorities to maximize plan quality.
+Unlike population-based approaches, our framework trains personalized agents for each patient using their planning Computed Tomography (CT) and augmented anatomies simulating anatomical changes (tumor progression and regression).
 This patient-specific approach leverages anatomical similarities along the treatment course, enabling effective plan adaptation.
 We implemented and compared two DRL algorithms, Deep Q-Network (DQN) and Proximal Policy Optimization (PPO), using dose-volume histograms (DVHs) as state representations and a -dimensional action space of priority adjustments.
-Evaluation on five HNC patients using actual replanning CT data showed that both DRL agents improved initial plan scores from to (DQN) and (PPO), surpassing the replans manually generated by a human planner ().
-Clinical validation confirms these improvements translate to better tumor coverage and OAR sparing across diverse anatomical changes.
+Evaluation on eight HNC patients using actual replanning CT data showed that both DRL agents improved initial plan scores from to (DQN) and (PPO), surpassing the replans manually generated by a human planner ().
+Further comparison of dosimetric endpoints confirms these improvements translate to better tumor coverage and OAR sparing across diverse anatomical changes.
 This work highlights the potential of DRL in addressing the geometric and dosimetric complexities of adaptive proton therapy, offering a promising solution for efficient offline adaptation and paving the way for online adaptive proton therapy.
@@
 However, the Bragg peak makes IMPT highly sensitive to interfractional anatomical changes, such as tumor progression or regression, weight loss, and edema, that alter tissue density along the beam path (Huiskes et al., 2023; Sonke et al., 2019).
-These anatomical changes can shift the Bragg peak, degrading tumor coverage and/or OAR sparing, and often necessitate one or multiple manual replanning sessions during treatment to maintain clinical quality.
+These anatomical changes can shift the Bragg peak, degrading tumor coverage and/or OAR sparing, and often necessitate one or multiple manual replanning sessions during the treatment course to maintain clinical quality.
 Addressing these anatomical changes currently requires manual intervention: a multi-step process involving patient reimaging, anatomical recontouring, and iterative, trial-and-error plan optimization.
-This manual adaptation workflow is notoriously labor-intensive and typically takes several days, delaying optimal care and straining limited clinical resources (Kim et al., 2018; Leeman et al., 2017; Li et al., 2020).
+This manual adaptation workflow is notoriously labor intensive and typically takes several days, delaying optimal care and straining limited clinical resources (Kim et al., 2018; Leeman et al., 2017; Li et al., 2020).
 From a machine learning (ML) perspective, automating treatment planning presents several challenges.
@@
 Specifically, RL holds the potential to automate the iterative decision-making process of priority adjustments in treatment planning (Eckardt et al., 2021; Ebrahimi and Lim, 2021; Moreau et al., 2021; Yang et al., 2024).
-Furthermore, RL’s inherent adaptability aligns with the vision of personalized online adaptive radiotherapy, where treatment plans can be dynamically adapted based on a patient’s tumor response and anatomical changes during treatment.
+Furthermore, RL’s inherent adaptability aligns with the vision of personalized online adaptive radiotherapy, where treatment plans can be dynamically adapted based on a patient’s tumor response and anatomical changes along the treatment course.
 To realize this potential for automated IMPT replanning, we develop a patient-specific deep reinforcement learning (DRL) framework for the dosimetrically challenging HNC.
@@
 Plan quality is quantified via a comprehensive -point scoring system that combines ProKnow standardized scoring criteria (Nelms et al., 2012) with institutional planning guidelines, and the reward is defined as the change in the plan quality score.
-Our framework trains personalized DRL agents using each patient’s anatomy captured by the initial planning CT (Computed Tomography) images and augmented anatomies simulating anatomical variations (tumor regression/progression).
+Our framework trains personalized DRL agents using each patient’s anatomy captured by the initial planning Computed Tomography (CT) images and augmented anatomies simulating anatomical variations.
 This patient-specific approach leverages the patient’s inherent anatomical consistency along the treatment course to optimize performance specifically for that individual rather than attempting to develop a one-size-fits-all solution.
@@
 Gao et al.
-(2024) adapted the hierarchical VTP approach for HNC photon therapy (whereas this work focuses on proton therapy), employing two DQN subnetworks for parameter selection and adjustment direction determination while representing states as plan quality scores rather than DVHs.
+(2024) adapted the hierarchical VTP approach for HNC photon therapy (whereas our work focuses on proton therapy), employing two DQN subnetworks for parameter selection and adjustment direction determination while representing states as plan quality scores rather than DVHs.
 To manage the high dimensionality arising in HNC planning, they manually fixed most of the potential planning parameters in their case based on dosimetrist expertise, allowing their VTP to learn adjustments for only unique priority parameters.
@@
 Their approach proposed dynamically reducing complexity by adjusting parameters only for the subset of structures currently exhibiting the lowest quality scores.
-However, this strategy may not accommodate cases where certain OARs must be strategically compromised to maintain target coverage—a common requirement in complex HNC planning scenarios.
+However, this strategy may not accommodate cases where certain OARs have to be strategically compromised to maintain target coverage—a common situation in complex HNC planning scenarios.
 Notably, all the aforementioned approaches rely on population-based models trained across diverse patient cohorts.
 While effective at capturing common anatomical patterns, such models face challenges adapting to patient-specific anatomical changes that necessitate treatment replanning during HNC’s fractionated radiotherapy (Caudell et al., 2017; Ma et al., 2022; Maniscalco et al., 2023; Visak et al., 2024).
-This limitation, stemming from the under-utilization of intra-patient anatomical consistency, restricts their ability to generate optimally adapted plans tailored to each patient’s unique progression.
+This limitation, stemming from the under-utilization of intra-patient anatomical consistency, restricts their ability to generate optimally adapted plans tailored to each patient’s unique anatomical features and evolving pattern.
 Motivated by these challenges, we introduce a novel patient-specific RL framework for IMPT replanning in HNC.
@@
 For the patient cohort in this study, and .
-The weighting parameters and serve as treatment planning priorities (TPPs) that balance competing clinical objectives: achieving prescribed dose coverage in the -th CTV while minimizing dose exposure to the -th OAR.
-These weights are the primary optimization variables manipulated by our DRL agents.
 and denote the dose influence matrices for the -th CTV and -th OAR.
 is the prescribed dose vector for the -th CTV.
-In practice, achieving a perfect homogeneous dose distribution within CTVs is not feasible; hot and cold spots are often unavoidable.
-Similarly, since radiation beams must often pass through some OARs to reach the CTVs, complete OAR sparing from radiation dose is unrealistic.
-Importantly, different OARs have varying dose tolerances based on their biological characteristics; as long as the dose to each OAR remains within clinically acceptable thresholds, the risk of severe toxicity is minimized.
-Hence, to reflect these clinical realities and achieve realistic plan optimization, in this study we implemented a DVH-based cost function incorporating multiple planning objectives.
-The planning objectives involved in our IMPT plan optimization for HNC are listed in Table 1, where represents the volume of a structure receiving a dose equal to or greater than , while denotes the minimum dose received by the hottest volume of a structure.
-The trade-offs between these competing objectives are controlled by objective-specific priorities (), which are traditionally set manually through iterative tuning.
-The DVH-based cost function (Eqn.
-2) was optimized using an in-house dose calculation engine, which serves as the environment for our RL-based planning optimization.
+The weighting parameters and serve as treatment planning priorities (TPPs) that balance competing clinical objectives: achieving prescribed dose coverage in the -th CTV while minimizing dose exposure to the -th OAR.
+While this weighted loss function is a standard proxy for the clinical objective, it does not directly optimize for the clinical constraints that human planners must satisfy (e.g., specific DVH criteria for various organs listed in Table 1).
+In practice, achieving perfect dose homogeneity is not feasible, and different OARs have varying dose tolerances based on their biological characteristics.
+These clinical realities mean that planners must manually and iteratively adjust the weighting parameters and re-solve the optimization problem until clinical requirements are met–a time-consuming and resource-intensive process.
+We automate this critical step by introducing a DRL agent to perform this hyperparameter search.
+The agent learns to directly optimize for clinical constraint satisfaction, using a reward signal derived from the dose plan’s adherence to clinical goals.
+The DRL agent interacts with a plan optimization engine that serves as the core of our RL environment, minimizing the cost function (Eqn.
+2) with agent-selected priority weights.
+This integration of dose calculation with RL-based planning optimization streamlines a major bottleneck in the radiotherapy workflow.
 3.2 Reinforcement Learning for Automatic Replanning
@@
 Action Space (): A discrete set of clinically constrained actions representing priority adjustments for different planning objectives, designed based on clinical experience.
+Adjustment details are provided in Appendix C.
 Transition Dynamics (): The transition model is governed by a treatment optimization engine, with capturing how priority adjustment actions affect the resulting dose distribution.
@@
 The final score was calculated as the sum of the scores across all these metrics, with higher scores indicating better plan quality.
+The plot of the scoring function for each dosimetric metric is presented in Appendix D.
 Discount Factor (): balances immediate dosimetric improvements with long-term overall plan quality.
@@
 To address this challenge, we implemented two state-of-the-art DRL algorithms that have demonstrated success in complex sequential decision-making tasks.
-Deep Q-Network (DQN)
+Deep Q-Network (DQN):
 In Q-learning, the objective is to learn the optimal action-value function , which represents the maximum expected return achievable by taking action in state and following the optimal policy thereafter (Sutton et al., 1998; Watkins and Dayan, 1992).
@@
 The DQN agent learns a policy by selecting actions with the highest Q-value for a given state (e.g., using an -greedy strategy to balance exploration and exploitation).
-Proximal Policy Optimization (PPO)
+Proximal Policy Optimization (PPO):
 While DQN learns Q-values of state-action pairs and derives policies through action selection mechanisms, PPO takes a more direct approach.
@@
 4.1 Cohort Selection
-We validated our approach through a retrospective study involving five HNC patients treated with IMPT at Hospital A.
-All patients underwent replanning during their treatment course due to substantial anatomical changes.
-A CT scan for treatment planning (referred to as pCT) was acquired weeks before the treatment course, and a second CT scan was obtained for replanning (rpCT) during the course when substantial anatomical changes were observed and warranted replanning.
-These five cases were deliberately selected for their varied tumor characteristics and spatial relationships to surrounding OARs, and all cases involved three CTVs, representing challenging scenarios for treatment planning.
-The primary CTV (CTV1) for all cases was prescribed to receive over fractions, while the prescription doses for the secondary (CTV2) and tertiary (CTV3) targets varied among cases, as detailed in Table 2.
-This variability in prescription doses poses a significant challenge for automated treatment planning and replanning in HNC, as population-based models often struggle to generalize across diverse clinical objectives, let alone the substantial variation in patient anatomy and tumor characteristics.
-Case ID | Volume on pCT () | Volume on rpCT () | No. of Fractions | Prescription (Gy RBE) | ||||||
-CTV1 | CTV2 | CTV3 | CTV1 | CTV2 | CTV3 | CTV1 | CTV2 | CTV3 | ||
-P1 | 35 | |||||||||
-P2 | 35 | |||||||||
-P3 | 35 | |||||||||
-P4 | 35 | |||||||||
-P5 | 35 |
+We validated our approach using a retrospective cohort of HNC patients treated with IMPT at the Emory Proton Therapy Center.
+Eight patients requiring replanning due to substantial anatomical changes were used for patient-specific RL training, while the remaining 10 patients served as training data for population-based RL baselines.
+For each patient, an initial planning CT (pCT) was acquired weeks before the treatment course, and a replanning CT (rpCT) was obtained when significant anatomical changes were observed and warranted replanning.
+Performance evaluation used all eight rpCT cases for both RL approaches.
+These cases were selected for their varied tumor characteristics and spatial relationships to surrounding OARs, all involving three CTVs representing challenging planning scenarios.
+While the primary CTV (CTV1) for all cases was prescribed over fractions, the prescription doses for the secondary (CTV2) and tertiary (CTV3) targets varied per case (Table 2).
+This prescription heterogeneity, combined with substantial inter-patient anatomical variation, poses a major challenge for automated treatment planning in HNC, particularly for population-based models struggling to generalize across heterogeneous clinical objectives.
+Patient demographics and staging are detailed in Table 6 (Appendix B).
+[…]
+CTV1 | CTV2 | CTV3 | CTV1 | CTV2 | CTV3 | CTV1 | CTV2 | CTV3 | CTV1 | CTV2 | CTV3 | ||
+P1 | 2 | ||||||||||||
+P2 | 1 | NA | NA | NA | |||||||||
+P3 | 1 | NA | NA | NA | |||||||||
+P4 | 2 | ||||||||||||
+P5 | 1 | NA | NA | NA | |||||||||
+P6 | 2 | ||||||||||||
+P7 | 2 | ||||||||||||
+P8 | 2 |
 4.2 Data Extraction
@@
 The dose matrix calculated from the pCT and its associated contours was used for training the DRL agents, while the matrix calculated from the rpCT and corresponding contours was used for testing.
-To train a patient-specific agent capable of adjusting the priorities among planning objectives to adapt the treatment plan to anatomical changes, we further generated augmented anatomies by simulating anatomical changes.
+To train a patient-specific agent capable of adapting treatment plans to anatomical changes, we generated augmented anatomies by simulating tumor variations.
 In this proof-of-concept study, we simulated only two variation scenarios, i.e., tumor progression and regression, by expanding the original CTV contours on the pCT by or shrinking them by , respectively.
@@
 In this section, we present the results of our experiments and evaluate the effectiveness of RL-based planners for automated IMPT replanning in HNC treatment.
-Case ID | Initial plan | Manual plan | DQN plan | PPO plan |
-P1 | 122.56 | 130.85 | 131.05 | 137.88 |
-P2 | 83.14 | 132.09 | 136.20 | 138.05 |
-P3 | 132.35 | 140.44 | 147.25 | 147.52 |
-P4 | 132.14 | 143.94 | 146.10 | 148.77 |
-P5 | 132.96 | 138.67 | 138.32 | 141.50 |
-mean std |
 5.1 Experimental Setup
@@
 Each agent was trained for episodes, with each episode consisting of up to priority adjustment steps or terminating early if the maximum plan quality score of was achieved.
-Algorithm 1 provides a detailed overview of this DRL-based replanning process, illustrating the integration of priority tuning and plan optimization within our experimental framework.
+Algorithm 1 in Appendix A provides a detailed overview of this DRL-based replanning process, illustrating the integration of priority tuning and plan optimization within our experimental framework.
+Details of network specifications and hyperparameters are provided in Appendix E.
 The trained agents were then tested on the patient’s new anatomy captured in the rpCT.
-Specifically, starting from an initial default priority set , the agents performed priority tuning within tuning steps, where each step involved selecting a tuning action based on the plan quality of the intermediate plan and was followed by inverse plan optimization.
-For comparison, a new treatment plan was also manually generated for each case on the new anatomy by a human planner through manual priority tuning.
+Specifically, starting from an initial default priority set , the agents performed priority tuning following the same episodic framework described above, where each step involved selecting a tuning action based on the intermediate plan’s quality, followed by inverse plan optimization.
+The number of tuning attempts (horizon length) in this process can be extended as needed, bounded only by computational cost.
+As a comparative baseline, a human planner also generated a new manual treatment plan for each case, employing manual priority tuning.
 The performance of the agents was assessed by comparing the plan quality scores and dosimetric metrics of the agent-generated plans with those of the manually created plans.
 To facilitate the interaction between the RL agents and the treatment planning process, we developed a custom OpenAI Gymnasium-compatible environment that simulates the treatment planning workflow.
-At each timestep, the environment processes the current beamlet weights along with the dose influence matrices and for each CTV and OAR .
-It then solves the DVH-based optimization problem using the priority set , modified by the agent-selected action , updates , and generates a new dose distribution , which is subsequently used to compute the plan score and reward via Eqn.
-1.
-Each action updates priority weights according to:
-where ensures weights remain within valid bounds. All experiments were conducted on a workstation equipped with an NVIDIA RTX Ada Generation GPU with of memory.
+At each time step, the environment receives the agent-selected action at and applies it to modify the current priority weights .
+Using these updated priorities along with the dose influence matrices and for each CTV and OAR , the environment solves the optimization problem (Eqn.
+2) via projected gradient descent (Fu et al., 2023; Ghobadi et al., 2012; Nocedal and Wright, 1999) to update the beamlet weights and generate a new dose distribution (Eqn.
+1).
+The resulting dose distribution is evaluated against clinical constraints (Table 1) to compute the plan score and reward (Appendix D).
+Each action updates priority weights according to
+where ensures weights remain within valid bounds.
+The details of the priority adjustment are included in the Appendix C.
+All experiments were conducted on a workstation equipped with an NVIDIA RTX Ada Generation GPU with of memory.
+Case ID | Initial | Manual | DQN-popn | DQN | PPO-popn | PPO |
+P1 | 122.56 | 130.85 | 130.40 | 131.05 | 137.01 | 137.88 |
+P2 | 83.14 | 132.09 | 135.33 | 136.20 | 135.64 | 138.05 |
+P3 | 132.35 | 140.44 | 144.17 | 147.25 | 147.31 | 147.52 |
+P4 | 132.14 | 143.94 | 148.38 | 146.10 | 146.38 | 148.77 |
+P5 | 132.96 | 138.67 | 138.41 | 138.32 | 138.30 | 141.50 |
+P6 | 125.65 | 133.03 | 128.87 | 135.24 | 135.18 | 135.18 |
+P7 | 108.67 | 132.53 | 126.27 | 141.03 | 136.51 | 141.35 |
+P8 | 128.75 | 139.03 | 124.31 | 141.53 | 137.06 | 141.78 |
+mean std |
 5.2 Plan Quality Comparisons
 The same -point scoring system used to calculate rewards during DRL training was employed to quantify the plan quality of treatment plans generated for each patient’s new anatomy captured in the rpCT.
-The resulting plan quality scores are presented in Table 3.
-The score for each dosimetric metric of the generated plans are summarized in Appendix A.
-As shown, both DQN and PPO agents effectively adjusted the priorities of planning objectives, significantly improving the plan quality compared to the initial plans generated using the default priority set.
-Notably, for every case, both agents achieved plans that were comparable to or better than those manually created by a human planner.
-PPO-generated plans consistently achieved the highest quality scores across all cases (mean: ), outperforming both manual plans (, via paired t-test) and DQN-generated plans ().
-Furthermore, the PPO-generated plans exhibited lower score variability (, vs.
-for manual plans and for DQN-generated plans), indicating enhanced consistency and reliability in adapting to anatomical changes.
+To provide comprehensive benchmarking, we evaluated our patient-specific approach against both manual planning and population-based RL baselines.
+The population-based agents were trained on the initial planning CTs of the remaining patients in our cohort using identical architecture, hyperparameters, and training setup, consistent with conventional clinical practice where models are trained on population data.
+[…]
+Overall, patient-specific agents matched or exceeded manual plans in all cases, with PPO achieving the highest score in out of patients.
+Beyond manual benchmarks, patient-specific agents significantly outperformed population-based agents in of the comparisons (across all patients and algorithms), underscoring the clear advantages of personalized adaptation over generalized models, especially when handling unique anatomical changes.
 5.3 Dosimetric Analysis
-The detailed dosimetric metrics of each structure for the manually generated plans and those generated by the DQN and PPO agents are presented in Table 4.
-Clinically acceptable target dose coverage was achieved in all the three plans for every case, with the coverage of primary CTV () no less than .
-In terms of dose homogeneity inside the primary CTV (CTV1), both RL agents reduced hot spots in three patients while maintaining excellent coverage.
-Most notably, in patient P4, DQN and PPO reduced the maximum dose within CTV1 by and , respectively, compared to the manual plan.
-PPO demonstrated superior OAR sparing in most patients, reducing mean dose to the larynx (LAR) ( (PPO) vs.
-(Manual)) and to the bilateral parotids (PARL and PARR) ( (PPO) vs.
-(Manual)), compared to manual replanning.
-The effectiveness of PPO was particularly evident in case P5, where it slightly compromised the dose coverage of the tertiary CTV (CTV3, ) to substantially reduce the dose to the right submandibular gland (SMGR), bringing its mean dose closer to the dose tolerance ( ) and resulting in a significantly higher plan score.
-In contrast, the manual and DQN-generated plans for this case substantially exceeded the dose tolerance of the SMGR, with mean doses of and , respectively.
-Figure 1 compares the DVH of the plans generated manually by a human planner against the automated plans generated using RL agents (DQN and PPO) for patients P4 and P5.
-DVH curves and dosimetric results for all cases are provided in Appendix A.
+[…]
 Structure | Metric | P1 | P2 | P3 | P4 | P5 | ||||||||||
@@
 Note: OAR abbreviations - BRS: Brainstem, SC: Spinal Cord, MAN: Mandible, LAR: Larynx, PHY: Pharynx, PARL/PARR: Left/Right Parotid, COCHL/COCHR: Left/Right Cochlea, SMLG/SMGR: Left/Right Submandibular Gland, ESO: Esophagus.
+5.4 Adaptation to Inter-fractional Changes
+To further assess the robustness of our approaches in handling rapidly evolving anatomies, we extended our comparison to include the second replanning CT for cases in our evaluation cohort that underwent multiple replanning sessions during the treatment course (P1, P4, P6, P7, P8).
+As shown in Table 5, patient-specific RL demonstrated superior adaptation capabilities, achieving better performance than population-based baselines in out of metric comparisons (DQN: cases; PPO: cases).
+Patient-specific DQN, in particular, achieved a higher mean score of compared to for the population-based DQN baseline, while showing markedly lower variability.
+Although population-based approaches showed efficacy in handling certain cases with common anatomical patterns (e.g., P4 with DQN), the patient-specific approach consistently demonstrated better adaptation to the unique inter-fractional changes of individual patients.
+Case ID | DQN-popn | DQN | PPO-popn | PPO |
+P1 | 124.95 | 133.33 | 126.78 | 133.64 |
+P4 | 146.23 | 143.13 | 145.11 | 145.88 |
+P6 | 132.55 | 135.69 | 135.16 | 136.54 |
+P7 | 115.27 | 140.57 | 135.21 | 141.11 |
+P8 | 119.92 | 139.42 | 130.26 | 138.92 |
+mean std |
 6 Discussion
-Our study demonstrates that RL-based automated IMPT replanning, particularly using PPO, consistently generates treatment plans of higher quality for HNC patients experiencing anatomical changes during the treatment course, compared to those manually generated by a human planner.
+Our study demonstrates that patient-specific RL-based automated IMPT replanning, particularly using PPO, consistently generates treatment plans of higher quality for HNC patients experiencing anatomical changes during the treatment course, compared to both manually generated plans and population-based RL approaches.
 These improvements have several important implications for adaptive proton therapy workflows.
 6.1 Clinical Significance
-The marked improvement in plan quality scores across all patients indicates that RL-based automated replanning can standardize high-quality treatment planning.
-This may lead to reduced treatment-related toxicities while maintaining effective tumor control.
+The marked improvement in plan quality scores across all patients indicates that patient-specific RL-based automated replanning can standardize high-quality treatment planning tailored to individual anatomical variations.
+This personalized approach may lead to reduced treatment-related toxicities while maintaining effective tumor control.
 For example, the reduction in parotid gland dose has been correlated with decreased risk of xerostomia, with studies suggesting that each reduction in mean dose translates to approximately reduction in xerostomia risk (Castelli et al., 2015; Chao et al., 2001).
 As shown in Table 4, patients P1 and P2 benefited from substantial parotid sparing (reductions of and , respectively), while patients P1 and P4 experienced notable laryngeal sparing (reductions of and ).
-Importantly, these dosimetric improvements were achieved without compromising target coverage, underscoring the ability of RL methods to address the geometric and dosimetric complexities in HNC treatment planning and balance complex trade-offs to achieve clinically acceptable treatment plans.
-The proposed RL-based automated replanning approach offers a promising solution for efficient offline adaptation and paves the way for online adaptive proton therapy.
+Importantly, these dosimetric improvements were achieved without compromising target coverage, underscoring the ability of patient-specific RL methods to address the geometric and dosimetric complexities unique to each patient’s HNC anatomy and balance complex trade-offs to achieve clinically acceptable treatment plans.
+The superior performance of patient-specific agents over population-based approaches ( comparisons in Table 3) demonstrates that individualized adaptation is crucial for handling the substantial inter-patient anatomical variability characteristic of HNC.
+The proposed patient-specific RL-based automated replanning approach offers a promising solution for efficient offline adaptation and paves the way for personalized online adaptive proton therapy.
 6.2 RL-based Planning and Patient-specific Benefits
-Both RL algorithms outperformed manual replanning, with PPO consistently delivering superior plans compared to DQN despite operating within the same discrete action space.
+Patient-specific RL algorithms outperformed both manual replanning and population-based counterparts, with PPO consistently delivering superior plans compared to DQN.
 This finding aligns with recent literature suggesting policy-based methods like PPO effectively navigate the complex reward landscapes common in radiotherapy planning (Wang and Chang, 2024).
-PPO’s strategy of performing small, incremental parameter updates helps prevent detrimental large policy shifts, a feature particularly well-suited to the radiotherapy optimization problem.
-Analysis of patient-specific outcomes relative to anatomical features (Tables 2 and 3) underscores this advantage.
-PPO consistently delivered the highest plan scores, outperforming manual replans across the cohort irrespective of initial target volumes (rpCT CTV1 range: ) or the extent of their volumetric changes between pCT and rpCT (e.g., P2 CTV3: expansion).
-Notably, it achieved the highest absolute score () for the patient with the smallest CTV1/CTV2 (P4), while providing clinically meaningful gains over manual plans for large-volume cases (PPO plan score vs.
-manual for P3: ; P5: ).
-Overall, the results strongly suggest that DRL-based automation offers significant dosimetric advantages for IMPT replanning.
-By consistently optimizing plans towards high scores, the DRL framework holds considerable promise for improving plan quality for HNC patients undergoing significant anatomical changes.
+PPO’s strategy of performing small, incremental parameter updates helps prevent detrimental large policy shifts, a feature particularly well-suited to patient-specific radiotherapy optimization where individual anatomical nuances require careful navigation.
+[…]
+[…]
+These findings collectively demonstrate that patient-specific DRL delivers significant dosimetric advantages over population-averaged approaches by learning and adapting to individual patient characteristics.
+By consistently delivering high-quality plans tailored to each patient’s evolving anatomy, this framework holds considerable promise for improving adaptive radiotherapy outcomes in HNC treatment.
 6.3 Limitations and Future Directions
 While our findings demonstrate the feasibility of deep reinforcement learning (DRL) for adaptive IMPT replanning, important limitations should be acknowledged alongside opportunities for future work.
-First, the retrospective evaluation was conducted on a limited cohort of five patients.
+First, the retrospective evaluation was conducted on a limited cohort of eight patients.
 However, this cohort is representative of challenging HNC cases with three CTVs of different prescription levels and common anatomical change patterns.
@@
 This project is supported by the National Institute of Health under Award Numbers
-R01DE033512 and R37CA272755.
-Appendix A Complete Dosimetric Results
+R01DE033512 and R37CA272755. We thank Dr. David S. Yu, Dr. Mark McDonald, and Dr. Ralph Weichselbaum for providing valuable clinical perspectives that informed this study.
+Appendix A Algorithm: Patient-Specific DRL for IMPT Replanning
+Appendix B Cohort Demographics and Staging
+Table 6 summarizes the demographic and clinical staging data for the eight patients included in this study.
+Patient ages range from to years.
+The cohort encompasses a range of disease stages, including stage II (), stage III (), stage IVa (), stage IVb (), and one unassigned case.
+Detailed TNM classifications are also reported, reflecting the heterogeneity of tumor and nodal involvement across the sample.
+Case ID | Age | Overall Stage | TNM Stage |
+P1 | 69 | IVb | T0 N3 M0 |
+P2 | 73 | Unassigned | T4 & T2 N2 M0 |
+P3 | 60 | II | T3 N2 M0 |
+P4 | 55 | III | T4 N0 M0 |
+P5 | 48 | II | T2 N2 M0 |
+P6 | 72 | IVa | T4 N0 M0 |
+P7 | 71 | III | T4 N0 M0 |
+P8 | 59 | III | T4 N1 M0 |
+Appendix C Priority Adjustments: Structure-Specific Modifications
+All replanning agents modify the treatment plan optimization priorities through discrete adjustments to weight parameters.
+The complete mapping is detailed in Table 7, organized by structure type and adjustment magnitude.
+This discrete action space enables clinically meaningful trade-off adjustments while maintaining valid priority ranges.
+For critical OARs (brainstem, spinal cord), larger adjustments () are available to prioritize their protection, while other structures have more moderate adjustments.
+Target volumes include both positive and negative adjustments to allow for both improved coverage and compromises when necessary for OAR sparing.
+Action Index | Structure | Parameter | Adjustment () |
+Target Volume Adjustments | |||
+0, 1 | CTV1 | ||
+2, 3 | CTV2 | ||
+4, 5 | CTV3 | ||
+18, 19 | CTV2 | ||
+20, 21 | CTV3 | ||
+Organ-at-Risk Adjustments | |||
+6 | Mandible (MAN) | ||
+7 | Brainstem (BRS) | ||
+8 | Spinal Cord (SC) | ||
+9, 10 | Parotids (PARR, PARL) | ||
+11 | Larynx (LAR) | ||
+12 | Pharynx (PHY) | ||
+13, 14 | Cochleas (COCHL, COCHR) | ||
+15, 16 | Submand. Glands (SMGL, SMGR) | ||
+17 | Esophagus (ESO) |
+Appendix D Plan Quality Metrics
+Plan quality was quantified using a comprehensive 150-point scoring system derived from standardized ProKnow criteria (Nelms et al., 2012) and institutional planning guidelines.
+This total score is the sum of individual component scores (detailed in Figure 2), which directly relate to the planning objectives outlined in Table 1.
+For reinforcement learning, the reward signal was defined as the change in this cumulative plan quality score between tuning steps.
+Appendix E Network Architectures & Training Parameters
+Both DRL agents process the DVHs using Convolutional Neural Networks (CNNs).
+This architectural choice is motivated by the ability of CNNs to learn meaningful spatial features and correlations from the structured DVH input, capturing complex dose-volume relationships across different anatomical structures.
+Common elements in both architectures include ReLU activation functions, batch normalization layers following convolutions, max-pooling for dimensionality reduction, and dropout layers for regularization.
+The DQN agent employs two convolutional layers followed by three fully connected layers to estimate Q-values for discrete actions.
+The DQN implementation features experience replay with prioritized sampling based on temporal difference errors, exponential epsilon-greedy decay, and a delayed target network updated every steps to stabilize learning.
+The PPO agent adopts a shared actor-critic architecture with a three-layer CNN backbone.
+The actor head outputs discrete priority adjustments via a categorical distribution, while the critic estimates state values.
+Training incorporates clipped policy updates (), generalized advantage estimation (GAE, ), value loss coefficient (), and entropy coefficient () to balance exploration and exploitation.
+Orthogonal initialization was applied to weights in the PPO network, with constant initialization for biases.
+Both algorithms use the Adam optimizer with learning rate , with PPO additionally employing gradient clipping (max norm ) to ensure training stability.
+Table 8 details the architectural specifications for both models.
+Table 9 highlights the key hyperparameters and optimization settings used in training.
+Component | DQN | PPO |
+Input Dimension | 15 × 100 (Structures × Dose Bins) | |
+Input Channels | 1 | 1 |
+Convolutional Layers | 2 | 3 |
+Kernel Sizes | (3,5) | (3,5) |
+Padding | (1,2) | (1,2) |
+Pooling | 2×2 MaxPool | 2×2 MaxPool |
+BatchNorm | Yes | Yes |
+Dropout Rate | 0.2 | 0.2 |
+Hidden Dim (CNN) | 32 → 64 | 32 → 64 → 64 |
+Fully Connected | 256 → 128 → 22 | 512 → 22 (Actor) |
+Output Activation | Linear | Softmax (Actor) |
+Parameter | DQN | PPO | ||
+Optimizer | Adam | Adam | ||
+Learning Rate | 1e-5 | 1e-5 | ||
+Minibatch Size | ||||
+Discount Factor () | ||||
+Loss Function | MSE |
+(Eqn. 3) | (Eqn. 5) | |||
+Experience Buffer | Prioritized Replay | On-policy Rollouts | ||
+Buffer Retention | Long-term | Episodic | ||
+Epsilon-Greedy Exploration | Yes (exponential decay) | No | ||
+Advantage Estimation | – | GAE () | ||
+Clipped Policy Update | – | |||
+Value Clip | – | |||
+Value Loss Coefficient () | – | |||
+Entropy Coefficient () | – | |||
+Gradient Clipping | Optional | (max norm) |
+Appendix F Complete Dosimetric Results
+[…]
 Delineated | Dosimetric Endpoint | Manual | DQN | PPO | ||||
@@
 CTV3 | 15 | 98.02 | 15.00 | 97.94 | 14.96 | 98.03 | 15.00 | |
-BRS | 21 | 2.87 | 21.00 | 2.82 | 21.00 | 20.04 | 21.00 | |
-SC | 21 | 13.85 | 21.00 | 14.04 | 21.00 | 22.85 | 21.00 | |
+BRS | 21 | 2.87 | 21.00 | 2.82 | 21.00 | 2.85 | 21.00 | |
+SC | 21 | 13.85 | 21.00 | 14.04 | 21.00 | 14.98 | 21.00 | |
 MAN | 5 | 0.22 | 5.00 | 0.09 | 5.00 | 0.09 | 5.00 | |
@@
 SMGL | 5 | 0.00 | 5.00 | 0.00 | 5.00 | 0.00 | 5.00 | |
-SMGR | 5 | 36.87 | 4.25 | 37.30 | 4.08 | 37.98 | 5.00 | |
+SMGR | 5 | 36.87 | 4.25 | 37.30 | 4.08 | 31.98 | 5.00 | |
 ESO | 5 | 1.36 | 5.00 | 1.38 | 5.00 | 1.38 | 5.00 | |
@@
 Cumulative Total | 150 | – | 138.67 | – | 138.32 | – | 141.50 |
+Delineated | Dosimetric Endpoint | Manual | DQN | PPO | ||||
+Structure | Metric | Score | Value | Score | Value | Score | Value | Score |
+CTV1 | 15 | 98.00 | 15.00 | 98.01 | 15.00 | 98.00 | 15.00 | |
+5 | 84.54 | 2.92 | 84.74 | 2.89 | 85.85 | 2.74 | ||
+CTV2 | 15 | 98.69 | 15.00 | 98.41 | 15.00 | 98.51 | 15.00 | |
+CTV3 | 15 | 98.46 | 15.00 | 98.15 | 15.00 | 98.39 | 15.00 | |
+BRS | 21 | 1.11 | 21.00 | 1.03 | 21.00 | 1.23 | 21.00 | |
+SC | 21 | 45.13 | 13.44 | 45.38 | 13.31 | 45.43 | 13.29 | |
+MAN | 5 | 0.00 | 5.00 | 0.00 | 5.00 | 0.00 | 5.00 | |
+LAR | 7 | 0.00 | 7.00 | 0.00 | 7.00 | 0.00 | 7.00 | |
+PHY | 7 | 0.00 | 7.00 | 0.00 | 7.00 | 0.00 | 7.00 | |
+PARL | 7 | 13.20 | 7.00 | 14.66 | 7.00 | 14.80 | 7.00 | |
+PARR | 7 | 13.67 | 7.00 | 14.13 | 7.00 | 14.31 | 7.00 | |
+COCHL | 5 | 0.00 | 5.00 | 0.00 | 5.00 | 0.00 | 5.00 | |
+COCHR | 5 | 0.00 | 5.00 | 0.00 | 5.00 | 0.00 | 5.00 | |
+SMGL | 5 | 36.19 | 4.53 | 34.94 | 5.00 | 34.97 | 5.00 | |
+SMGR | 5 | 39.59 | 3.16 | 35.01 | 5.00 | 34.96 | 5.00 | |
+ESO | 5 | 52.75 | 0.00 | 49.92 | 0.04 | 49.69 | 0.15 | |
+Cumulative Total | 150 | – | 133.03 | – | 135.24 | – | 135.18 |
+Delineated | Dosimetric Endpoint | Manual | DQN | PPO | ||||
+Structure | Metric | Score | Value | Score | Value | Score | Value | Score |
+CTV1 | 15 | 98.01 | 15.00 | 97.99 | 14.99 | 97.99 | 14.99 | |
+5 | 78.64 | 3.77 | 82.44 | 3.22 | 84.30 | 2.96 | ||
+CTV2 | 15 | 97.63 | 14.76 | 96.98 | 14.32 | 96.86 | 14.24 | |
+CTV3 | 15 | 98.08 | 15.00 | 97.16 | 14.44 | 97.30 | 14.53 | |
+BRS | 21 | 23.07 | 21.00 | 19.85 | 21.00 | 20.23 | 21.00 | |
+SC | 21 | 34.99 | 18.50 | 30.29 | 20.85 | 30.27 | 20.87 | |
+MAN | 5 | 0.03 | 5.00 | 0.07 | 5.00 | 0.21 | 5.00 | |
+LAR | 7 | 42.36 | 6.05 | 41.64 | 6.34 | 40.61 | 6.76 | |
+PHY | 7 | 50.47 | 6.81 | 48.31 | 7.00 | 48.50 | 7.00 | |
+PARL | 7 | 35.68 | 0.00 | 25.86 | 7.00 | 25.61 | 7.00 | |
+PARR | 7 | 23.26 | 7.00 | 23.10 | 7.00 | 25.32 | 7.00 | |
+COCHL | 5 | 20.38 | 5.00 | 18.11 | 5.00 | 18.43 | 5.00 | |
+COCHR | 5 | 7.54 | 5.00 | 5.31 | 5.00 | 5.60 | 5.00 | |
+SMGL | 5 | 67.71 | 5.00 | 66.30 | 0.00 | 66.60 | 0.00 | |
+SMGR | 5 | 35.91 | 4.63 | 35.35 | 4.86 | 34.84 | 5.00 | |
+ESO | 5 | 14.48 | 5.00 | 12.24 | 5.00 | 12.65 | 5.00 | |
+Cumulative Total | 150 | – | 132.53 | – | 141.03 | – | 141.35 |
+Delineated | Dosimetric Endpoint | Manual | DQN | PPO | ||||
+Structure | Metric | Score | Value | Score | Value | Score | Value | Score |
+CTV1 | 15 | 97.99 | 14.99 | 97.99 | 14.99 | 98.02 | 15.00 | |
+5 | 82.89 | 3.16 | 84.47 | 2.93 | 83.44 | 3.08 | ||
+CTV2 | 15 | 97.86 | 14.91 | 98.06 | 15.00 | 97.99 | 14.99 | |
+CTV3 | 15 | 97.77 | 14.85 | 97.51 | 14.68 | 97.51 | 14.68 | |
+BRS | 21 | 15.29 | 21.00 | 15.94 | 21.00 | 15.74 | 21.00 | |
+SC | 21 | 31.87 | 20.06 | 32.14 | 19.93 | 31.85 | 20.07 | |
+MAN | 5 | 0.46 | 5.00 | 0.52 | 5.00 | 0.46 | 5.00 | |
+LAR | 7 | 41.84 | 6.26 | 38.99 | 7.00 | 39.90 | 7.00 | |
+PHY | 7 | 0.00 | 7.00 | 0.00 | 7.00 | 0.00 | 7.00 | |
+PARL | 7 | 28.00 | 5.00 | 25.98 | 7.00 | 25.86 | 7.00 | |
+PARR | 7 | 17.32 | 7.00 | 17.20 | 7.00 | 17.31 | 7.00 | |
+COCHL | 5 | 4.15 | 5.00 | 4.40 | 5.00 | 4.37 | 5.00 | |
+COCHR | 5 | 0.00 | 5.00 | 0.00 | 5.00 | 0.00 | 5.00 | |
+SMGL | 5 | 69.03 | 0.00 | 68.98 | 0.00 | 68.77 | 0.00 | |
+SMGR | 5 | 35.51 | 4.80 | 34.32 | 5.00 | 35.11 | 4.96 | |
+ESO | 5 | 6.37 | 5.00 | 6.53 | 5.00 | 6.46 | 5.00 | |
+Cumulative Total | 150 | – | 139.03 | – | 141.53 | – | 141.78 |
+Structure | Metric | P6 | P7 | P8 | ||||||
+M | Q | P | M | Q | P | M | Q | P | ||
+CTV1 | 98.00 | 98.01 | 98.00 | 97.99 | 97.99 | 98.00 | 98.01 | 97.99 | 97.99 | |
+84.54 | 84.74 | 85.85 | 82.89 | 84.47 | 83.44 | 78.64 | 82.44 | 84.30 | ||
+CTV2 | 98.69 | 98.41 | 98.51 | 97.86 | 98.06 | 97.99 | 97.63 | 96.98 | 96.86 | |
+CTV3 | 98.46 | 98.15 | 98.39 | 97.77 | 97.51 | 97.51 | 98.08 | 97.16 | 97.30 | |
+BRS | 1.11 | 1.03 | 1.23 | 15.29 | 15.94 | 15.74 | 23.07 | 19.85 | 20.23 | |
+SC | 45.13 | 45.38 | 45.43 | 31.87 | 32.14 | 31.85 | 34.99 | 30.29 | 30.27 | |
+MAN | 0.00 | 0.00 | 0.00 | 0.46 | 0.52 | 0.46 | 0.03 | 0.07 | 0.21 | |
+LAR | 0.00 | 0.00 | 0.00 | 41.84 | 38.99 | 39.90 | 42.36 | 41.64 | 40.61 | |
+PHY | 0.00 | 0.00 | 0.00 | 0.00 | 0.00 | 0.00 | 50.47 | 48.31 | 48.50 | |
+PARL | 13.20 | 14.66 | 14.80 | 28.00 | 25.98 | 25.86 | 35.68 | 25.86 | 25.61 | |
+PARR | 13.67 | 14.13 | 14.31 | 17.32 | 17.20 | 17.31 | 23.26 | 23.10 | 25.32 | |
+COCHL | 0.00 | 0.00 | 0.00 | 4.15 | 4.40 | 4.37 | 20.38 | 18.11 | 18.43 | |
+COCHR | 0.00 | 0.00 | 0.00 | 0.00 | 0.00 | 0.00 | 7.54 | 5.31 | 5.60 | |
+SMGL | 36.19 | 34.94 | 34.97 | 69.03 | 68.98 | 68.77 | 67.71 | 66.30 | 66.60 | |
+SMGR | 39.59 | 35.01 | 34.96 | 35.51 | 34.32 | 35.11 | 35.91 | 35.35 | 34.84 | |
+ESO | 52.57 | 49.92 | 49.69 | 6.37 | 6.53 | 6.46 | 14.48 | 12.24 | 12.65 | |
+Plan Score (max=150): | 133.03 | 135.24 | 135.18 | 132.53 | 141.03 | 141.35 | 139.03 | 141.53 | 141.78 |
+Note: OAR abbreviations - BRS: Brainstem, SC: Spinal Cord, MAN: Mandible, LAR: Larynx, PHY: Pharynx, PARL/PARR: Left/Right Parotid, COCHL/COCHR: Left/Right Cochlea, SMLG/SMGR: Left/Right Submandibular Gland, ESO: Esophagus.